OpenStudy (anonymous):

Find all the zeroes of the equation. x^3 – x^2 = –3x – 3

hartnn (hartnn):

you sure about the Q ? because that has 2 imaginary root, and one irrational root..

hartnn (hartnn):

*2 complex roots

OpenStudy (anonymous):

yes I'm sure about the question. I seriously don't know what i'm doing or where to begin

hartnn (hartnn):

then that makes 2 of us...

OpenStudy (hba):

I think there's a mistake in the equation.

OpenStudy (anonymous):

What type of class is it? If it's fairly general maths, then they're not going to ask you to find a root that is actually irrational. To do that you'd need to use an algorithm to approximate it, or something like that. And it doesn't sound like that's what you'd be doing.

hartnn (hartnn):

i so wanted that to be +3 instead of -3

OpenStudy (hba):

The question should be : x^3 – x^2 = –3x + 3

OpenStudy (hba):

Do you have options ?

OpenStudy (anonymous):

I'm in algebra 2. and I copied the equation from my homework so there is no mistake

OpenStudy (hba):

@Jadelynn Show me the options please.

OpenStudy (anonymous):

the right hand side could also be 3x - 3. And your lecturer might have written it incorrectly. So that's where the mistake is.

OpenStudy (hba):

@Jadelynn Options ?

OpenStudy (anonymous):

hold on Imma find a way to show the options

OpenStudy (hba):

These are the options : A.i√3, –i√3, –1 B.i√3, –i √3, 1 C.√3, –√3, –1 D.i√3, 1 Right ?

hartnn (hartnn):

aah, so u have options! that simplifies it.

OpenStudy (anonymous):

@hba How'd you do that? and yea those are the options

OpenStudy (hba):

If the equation is x^3 - x^2 = -3x - 3 then none of your options are appropriate.

hartnn (hartnn):

because we have seen this question earlier

OpenStudy (hba):

Yeah lol and every time the post the wrong question.

hartnn (hartnn):

yes, every damn time....must be printing mistake...

hartnn (hartnn):

and correct Question is x^3 – x^2 = –3x + 3

hartnn (hartnn):

and correct answer is ...

OpenStudy (hba):

The question is x^3 – x^2 = –3x + 3 Now rewrite this equation as x^3-x^2+3x-3=0

OpenStudy (anonymous):

okay?

OpenStudy (hba):

Now Consider one of the roots to be 1 or -1 and then use polynomial division. and find the other roots.

hartnn (hartnn):

or, u can factor

OpenStudy (anonymous):

(x^3-x^2)+3x-3=0 x^2(x-1) + 3(x-1) = 0 (x-1) (x^2+3) =0

OpenStudy (anonymous):

nw x - 1 = 0 x = 1 or x^2+3 =0 x^2 = -3 x = sqrt(-3) =3i

OpenStudy (anonymous):

*really confused*

OpenStudy (anonymous):

\[x^3 - x^2 = -3x + 3 \] factor both sides, to get: \[x^2(x-1) = -3(x-1) \] move right hand side over to the left: \[x^2(x-1) + 3(x-1) = 0 \] there's a common factor of x-1, so: \[(x^2+3)(x-1) = 0\]

OpenStudy (hba):

@Jadelynn Did you understand the factoring part ?

OpenStudy (anonymous):

a little bit

OpenStudy (hba):

@Jadelynn Do you want an explanation on how to factorize ? Or should i continue ?

OpenStudy (anonymous):

um you can continue

OpenStudy (hba):

 Now solve this.

OpenStudy (hba):

@Jadelynn Solve for x

OpenStudy (hba):

So what did you get ?

OpenStudy (anonymous):

is it C?

OpenStudy (hba):

No, What answers did you get ?

OpenStudy (hba):

Do you know the complex root is \[\sqrt{-1}=i\]

OpenStudy (hba):

@Jadelynn ?

OpenStudy (anonymous):

@hba kinda

OpenStudy (hba):

So one of them is x=1 and \[x^2=-3\] \[x=\sqrt{-3}\] \[x=\sqrt{3}*\sqrt{-1},x=\sqrt{3}i \ or \ -\sqrt{3}i \]

OpenStudy (anonymous):

sorry my computer is severely lagging but thanks for helping

OpenStudy (hba):

No problem but you need to learn complex roots and factorizing Do you want me to teach you some questions ?